### v1

@@ A.1.2 Experimental setup
Generating data splits. There are three cell sets (i.e., groups of cells with the same cell state label) that consist of cells from Day 0.0 in our post-clustering version of the Waddington-OT dataset. We refer to these cell sets as “start clusters” because all initial cell states are sampled from one of these cell sets. Since the choice of start cluster can influence the likelihood of a cell’s trajectory reaching certain terminal fates, we split our cellular trajectories into train, validation, and test sets based on their start cluster. This cell-centric data split allows us to evaluate how well a model can generalize to different distributions of trajectories. Start cluster #1 is in the train set, start cluster #3 is in the validation set, and start cluster #2 is in the test set.
@@ A.2 Patient Lab Tests
A.2.1 Constructing routine lab test trajectories
@@
We leverage two publicly available medical datasets: eICU Pollard et al. (2018) and MIMIC-IV Johnson et al. (2024a; 2023); Goldberger et al. (2000). Both datasets are under the PhysioNet Credentialed Health Data License 1.5.0 PhysioNet . The retrieval process includes registering as a credentialed user on PhysioNet, completing the CITI “Data or Specimens Only Research” training, and signing the necessary data use agreements.
@@
Processing patient datasets. We process each dataset (i.e., eICU, MIMIC-IV) separately with the following steps. First, we extract the routine lab tests only (annotation available only in MIMIC-IV) and the most commonly ordered lab tests (i.e., lab tests that appear in at least 80% of patients). Next, we keep patients for whom we have at least one of each lab test. If there are multiple measurements of a lab test at the same time step (i.e., year, month, date, hour, minute, and seconds), we take the mean of its values. We extract patients with more than one visit (or time step).
@@
We define patients’ conditions as medical codes, specifically International Classification of Diseases (ICD), of their diagnosis. Both eICU and MIMIC-IV use ICD-9 and ICD-10 codes. We extract the medical codes and their timestamps (multiple medical codes at a single time step is possible). Since the timestamps of diagnostic codes and lab tests are not necessarily the same (and there are fewer entries of diagnostic codes than lab orders), we merge them with a tolerance range of 12 hours (eICU) or two days (MIMIC-IV). We obtain (frozen) condition embeddings $\mathbf{z}_{s}\in\mathbb{R}^{128}$ (retrieved on December 22, 2024) from an embedding model that has been pretrained on a clinical knowledge graph Johnson et al. (2024b). The clinical knowledge graph is constructed by integrating six existing databases of clinical vocabularies used in electronic health records: International Classification of Diseases (ICD), Anatomical Therapeutic Chemical (ATC) Classification, Systemized Nomenclature of Medicine - Clinical Terms (SNOMED CT), Current Procedural Terminology (CPT), Logical Observation Identifiers Names and Codes (LOINC), and phecodes Johnson et al. (2024b).
A.2.2 Generating data splits
@@
We generate a standard patient-centric random split for benchmarking model performance, and a series of increasingly challenging data splits via SPECTRA Ektefaie et al. (2024) to evaluate model generalizability.
@@
Constructing data splits to evaluate model generalizability. SPECTRA Ektefaie et al. (2024) creates a series of splits with decreasing cross-split overlap or similarity between the train and test sets. By training and testing models on these splits, we can assess model performance as a function of cross-split overlap. SPECTRA refers to this relationship as the spectral performance curve, which provides insight into how well a model generalizes to less similar data. When a new dataset split is encountered, it can be plotted as a point on this curve. The area under the spectral performance curve (AUSPC) serves as a metric of model generalizability and enables comparisons across models.
@@
To generate a split with SPECTRA, a similarity definition and a SPECTRA parameter (SP) value between 0 and 1 are required. SP controls the level of cross-split overlap: values closer to 0 create splits resembling classical random splits, while values closer to 1 produce stricter splits with minimal or no overlap between train and test sets. For example, at an input of 1, no similar samples are shared between the train and test sets.
@@
For eICU and MIMIC-IV, we define two patients as similar if: (1) they are of the same gender, (2) they are born in the same decade, and (3) they share at least one ICD-9 or ICD-10 category. We exclude ICD-9 and ICD-10 codes that are present in more than 50% of patients to avoid overly generic features. SPECTRA systematically prunes similar patients to produce splits. For this study, we generate 20 splits with SP values that are evenly spaced between 0 and 1. Given a train and test set, cross-split overlap is defined as the proportion of samples in the train set that are similar to at least one sample in the test set.
A.2.3 Constructing cohorts of patients with type 1 diabetes mellitus
@@
To define a type 1 diabetes mellitus (T1D) patient cohort in eICU and MIMIC-IV, we identify patients with T1D and matched healthy individuals. A patient has T1D if the ICD-10 code E10 (or the equivalent ICD-9 code 250) is present in the electronic health records. Matched healthy patients are defined by three criteria. First, the patient must not contain any of the following ICD-10 (and ICD-9 equivalent) codes: E11, E13, E12, E08, E09, R73, and O24. An initial healthy patient cohort is constructed using these filtering codes. Next, we identify frequently co-occurring ICD codes between the initial set of patients and patients with T1D to filter out generic ICD codes (threshold $=20$). Finally, healthy patients are matched with a T1D patient if: they are of the same gender, they are born in the same decade, and they share at least 50% of ICD codes.
Appendix B Further Implementation Details
@@
We provide code and instructions in our GitHub repository to implement Clef, baselines, and ablations: https://github.com/mims-harvard/CLEF. For the implementation of baselines, we followed the authors’ recommendations on model design and hyperparameter selection from the original publications.
B.1 Hyperparameter Sweep
@@
For all models trained from scratch, the selection of hyperparameters are: dropout rate $\in[0.3,0.4,0.5,0.6]$, learning rate $\in[0.001,0.0001,0.00001]$, and number of layers (or blocks in xLSTM) $\in[4,8]$. As the number of heads must be divisible by the number of features, the number of heads for eICU (18 lab tests) $\in[2,3,6,9]$ and for others $\in[4,8]$. For xLSTM, additional hyperparameters are: 1D-convolution kernel size $\in[4,5,6]$ and QVK projection layer block size $\in[4,8]$.
B.2 Best Hyperparameters
@@
MIMIC-IV dataset. The best hyperparameters for models trained on the MIMIC-IV dataset are: dropout rate $=0.6$, learning rate $=0.0001$, number of layers (or blocks in xLSTM) $=8$, and number of heads $=4$. For xLSTM models, 1D-convolution kernel size $=4$ and QVK projection layer block size $=4$. For Clef models, the number of FNN in the concept encoder $=1$ (Appendix Figures 8-9).
@@
eICU dataset. The best hyperparameters for models trained on the eICU dataset are: dropout rate $=0.6$, learning rate $=0.0001$, number of layers (or blocks in xLSTM) $=8$, and number of heads $=6$. For xLSTM models, the number of heads $=2$, 1D-convolution kernel size $=4$ and QVK projection layer block size $=4$. For Clef models, the number of FNN in the concept encoder $=1$ (Appendix Figures 8-9).
@@
Waddington-OT (WOT) dataset. The best hyperparameters for models trained on the WOT dataset are: dropout rate $=0.6$, learning rate $=0.00001$, number of layers (or blocks in xLSTM) $=4$, number of heads $=8$. For xLSTM models, 1D-convolution kernel size $=4$ and QVK projection layer block size $=8$. For Clef models, the number of FNN in the concept encoder $=0$ (Appendix Figures 8-9).
Appendix C Additional Figures and Tables
@@
Figure 8: Benchmarking the performance of Clef, baselines, and ablation models on (a) immediate and (b) delayed sequence editing. Performance is measured by MAE. The models are trained using a standard cell- or patient-centric random split. Not shown for visualization purposes are the performances of VAR models on eICU and MIMIC-IV datasets: on immediate sequence editing, MAE for eICU and MIMIC-IV are $55982.74$ and $886.05$, respectively; on delayed sequence editing, MAE for eICU and MIMIC-IV are $3.02\times 10^{39}$ and $8.62\times 10^{23}$, respectively.
@@
Figure 9: Benchmarking the performance of Clef, baselines, and ablation models on (a) immediate and (b) delayed sequence editing. Performance is measured by RMSE. The models are trained using a standard cell- or patient-centric random split. Not shown for visualization purposes are the performances of VAR models on eICU and MIMIC-IV datasets: on immediate sequence editing, MAE for eICU and MIMIC-IV are $135003.67$ and $1793.23$, respectively; on delayed sequence editing, MAE for eICU and MIMIC-IV are $5.84\times 10^{39}$ and $1.59\times 10^{24}$, respectively.
@@
Figure 10: Cross-split overlap (CSO) as a function of SPECTRA parameter (SP) for eICU and MIMIC-IV datasets. CSO is defined as the number of samples in the test set that are similar to at least one sample in the train set. SP is an internal parameter used by SPECTRA to control the CSO of generated data splits. CSO decreases as SP increases.
@@
Table 2: Generalizability of Clef, baselines, and ablations on eICU and MIMIC-IV datasets in immediate and delayed sequencing. Performance is measured by the area under the spectral performance curve (AUSPC) for MAE (Appendix Figure 11) or RMSE (Appendix Figure 12). Smaller AUSPC values indicate better performance.
@@
Clef-Transformer (FFN = 0) $15.16\pm 1.09$ $32.95\pm 2.47$ $14.36\pm 1.07$ $34.27\pm 2.12$ $32.79\pm 1.41$ $57.76\pm 3.39$ $35.65\pm 1.73$ $65.10\pm 4.43$
Clef-Transformer (FFN = 1) $10.99\pm 0.31$ $27.57\pm 0.27$ $9.25\pm 0.60$ $27.69\pm 0.22$ $21.35\pm 3.16$ $36.92\pm 5.46$ $23.83\pm 3.26$ $44.11\pm 5.83$
@@
Clef-xLSTM (FFN = 0) $16.73\pm 2.16$ $35.43\pm 6.01$ $15.32\pm 2.10$ $34.68\pm 7.09$ $32.06\pm 1.13$ $53.42\pm 2.18$ $33.88\pm 1.98$ $57.73\pm 3.63$
Clef-xLSTM (FFN = 1) $11.35\pm 0.11$ $28.09\pm 0.08$ $9.04\pm 0.18$ $26.21\pm 0.48$ $21.04\pm 2.32$ $37.50\pm 4.60$ $22.63\pm 2.61$ $42.12\pm 5.03$
@@
Clef-MOMENT (FFN = 0) $47.69\pm 0.33$ $82.18\pm 0.34$ $40.10\pm 0.44$ $72.70\pm 0.46$ $44.01\pm 0.35$ $73.83\pm 0.63$ $46.88\pm 0.38$ $81.20\pm 1.27$
Clef-MOMENT (FFN = 1) $47.56\pm 1.60$ $82.81\pm 2.88$ $39.91\pm 1.65$ $72.54\pm 3.20$ $42.92\pm 0.52$ $70.72\pm 1.96$ $45.75\pm 0.65$ $77.35\pm 2.77$
@@
Figure 11: Generalizability of Clef, baselines, and ablation models on (a) eICU and (b) MIMIC-IV patient datasets in immediate and delayed sequence editing. Performance is measured by MAE. As the SPECTRA parameter increases, the train/test split similarity decreases (Appendix Figure 10). The area under the spectral performance curve (AUSPC) evaluation is in Appendix Table 2.
@@
Figure 12: Generalizability of Clef, baselines, and ablation models on (a) eICU and (b) MIMIC-IV patient datasets in immediate and delayed sequence editing. Performance is measured by RMSE. As the SPECTRA parameter increases, the train/test split similarity decreases (Appendix Figure 10). The area under the spectral performance curve (AUSPC) evaluation is in Appendix Table 2.
@@
Figure 13: Benchmarking the performance of Clef, baselines, and ablation models on zero-shot (a) immediate and (b) delayed counterfactual generation of cellular developmental trajectories. Performance is measured by MAE (top row) and RMSE (bottom row).

### v2

@@ C.1.2 Experimental setup
Generating data splits. There are three cell sets (i.e., groups of cells with the same cell state label) that consist of cells from Day 0.0 in our post-clustering version of the Waddington-OT dataset. We refer to these cell sets as “start clusters" because all initial cell states are sampled from one of these cell sets. Since the choice of start cluster can influence the likelihood of a cell’s trajectory reaching certain terminal fates, we split our cellular trajectories into train, validation, and test sets based on their start cluster. This cell-centric data split allows us to evaluate how well a model can generalize to different distributions of trajectories. Start cluster #1 is in the train set, start cluster #3 is in the validation set, and start cluster #2 is in the test set.
@@ C.2 Patient Lab Tests
C.2.1 Constructing routine lab test trajectories
@@
We leverage two publicly available medical datasets: eICU [pollard2018eicu] and MIMIC-IV [johnson2024mimic, johnson2023mimic, goldberger2000physiobank]. Both datasets are under the PhysioNet Credentialed Health Data License 1.5.0 [mimiciv_license]. The retrieval process includes registering as a credentialed user on PhysioNet, completing the CITI “Data or Specimens Only Research" training, and signing the necessary data use agreements.
@@
We process each dataset (i.e., eICU, MIMIC-IV) separately with the following steps. First, we extract the routine lab tests only (annotation available only in MIMIC-IV) and the most commonly ordered lab tests (i.e., lab tests that appear in at least 80% of patients). Next, we keep patients for whom we have at least one of each lab test. If there are multiple measurements of a lab test at the same time step (i.e., year, month, date, hour, minute, and seconds), we take the mean of its values. We extract patients with more than one visit (or time step).
@@
We define patients’ conditions as medical codes, specifically International Classification of Diseases (ICD), of their diagnosis. Both eICU and MIMIC-IV use ICD-9 and ICD-10 codes. We extract the medical codes and their timestamps (multiple medical codes at a single time step is possible). Since the timestamps of diagnostic codes and lab tests are not necessarily the same (and there are fewer entries of diagnostic codes than lab orders), we merge them with a tolerance range of 12 hours (eICU) or two days (MIMIC-IV). We obtain (frozen) condition embeddings $\mathbf{z}_{s}\in\mathbb{R}^{128}$ (retrieved on December 22, 2024) from an embedding model that has been pretrained on a clinical knowledge graph [johnson2024unified]. The clinical knowledge graph is constructed by integrating six existing databases of clinical vocabularies used in electronic health records: International Classification of Diseases (ICD), Anatomical Therapeutic Chemical (ATC) Classification, Systemized Nomenclature of Medicine - Clinical Terms (SNOMED CT), Current Procedural Terminology (CPT), Logical Observation Identifiers Names and Codes (LOINC), and phecodes [johnson2024unified].
C.2.2 Generating data splits
@@
We generate a standard patient-centric random split for benchmarking model performance (R1-R2), and a series of increasingly challenging data splits via SPECTRA [ektefaie2024evaluating] to evaluate model generalizability (R3). We describe in detail the process of constructing SPECTRA data splits:
@@
SPECTRA [ektefaie2024evaluating] creates a series of splits with decreasing cross-split overlap or similarity between the train and test sets. By training and testing models on these splits, we can assess model performance as a function of cross-split overlap (Appendix Figure 10). SPECTRA refers to this relationship as the spectral performance curve, which provides insight into how well a model generalizes to less similar data. When a new dataset split is encountered, it can be plotted as a point on this curve. The area under the spectral performance curve (AUSPC) serves as a metric of model generalizability and enables comparisons across models (Appendix Table 2).
@@
To generate a split with SPECTRA, a similarity definition and a SPECTRA parameter (SP) value between 0 and 1 are required. SP controls the level of cross-split overlap (Appendix Figure 10): values closer to 0 create splits resembling classical random splits, while values closer to 1 produce stricter splits with minimal or no overlap between train and test sets. For example, at an input of 1, no similar samples are shared between the train and test sets.
@@
For eICU and MIMIC-IV, we define two patients as similar if: (1) they are of the same gender, (2) they are born in the same decade, and (3) they share at least one ICD-9 or ICD-10 category. We exclude ICD-9 and ICD-10 codes that are present in more than 50% of patients to avoid overly generic features. SPECTRA systematically prunes similar patients to produce splits. For this study, we generate 20 splits with SP values that are evenly spaced between 0 and 1 (Appendix Figure 10). Given a train and test set, cross-split overlap is defined as the proportion of samples in the train set that are similar to at least one sample in the test set (Appendix Figure 10).
C.2.3 Constructing cohorts of patients with type 1 diabetes mellitus
We conduct case studies on two independent cohorts of patients with type 1 diabetes mellitus (T1D), a chronic autoimmune disease in which the immune system attacks insulin-producing cells in the pancreas [quattrin2023type]. From our processed eICU and MIMIC-IV datasets, we construct two cohorts of T1D patients and matched healthy individuals.
@@ C.2.3 Constructing cohorts of patients with type 1 diabetes mellitus
Procedure. To define a type 1 diabetes mellitus (T1D) patient cohort in eICU and MIMIC-IV, we identify patients with T1D and matched healthy individuals. A patient has T1D if the ICD-10 code E10 (or the equivalent ICD-9 code 250) is present in the electronic health records. Matched healthy patients are defined by three criteria. First, the patient must not contain any of the following ICD-10 (and ICD-9 equivalent) codes: E11, E13, E12, E08, E09, R73, and O24. An initial healthy patient cohort is constructed using these filtering codes. Next, we identify frequently co-occurring ICD codes between the initial set of patients and patients with T1D to filter out generic ICD codes (threshold $=20$). Finally, healthy patients are matched with a T1D patient if: they are of the same gender, they are born in the same decade, and they share at least 50% of ICD codes.
Data statistics. eICU-T1D contains $59$ T1D patients and $579$ matched healthy controls, while MIMIC-IV-T1D includes $25$ T1D patients and $226$ matched healthy controls.
C.2.4 Experimental setup for type 1 diabetes mellitus case study
We evaluate Clef’s ability to simulate counterfactual patient trajectories through temporal concept intervention. This is analogous to intervening on concept bottleneck models by editing concept values and propagating the changes to the final prediction [koh2020concept]. Such a capability is particularly valuable when condition tokens are insufficient, such as when prescribing medication dosage. Editing concept values allow users, such as clinicians, to simulate potential trajectories as a result of the precise edits.
We conduct case studies on two independent cohorts of patients with type 1 diabetes mellitus (T1D) (Appendix C.2.4). To generate counterfactual sequences, we modify specific values in temporal concept $\mathbf{c}$, such as glucose levels, and allow Clef to simulate future trajectories of length $T=10$. We then compare these counterfactual trajectories (i.e., Clef-generated patients) against observed sequences from matched healthy individuals, other healthy individuals, and other T1D patients. Our hypothesis is that clinically meaningful edits will produce “healthier" (i.e., more similar to healthy patients) or “sicker" (i.e., more similar to other T1D patients) trajectories.
C.3 Synthetic tumor growth trajectories
The tumor growth simulation model [geng2017prediction] produces trajectories of tumor volume (i.e., one-dimensional outcome) after cancer diagnosis. There are two binary treatments (i.e., radiotherapy $\mathbf{A}^{r}_{t}$ and chemotherapy $\mathbf{A}^{c}_{t}$) at time $t$, and the possible treatments are: {($\mathbf{A}^{c}_{t}=0,\mathbf{A}^{r}_{t}=0$), ($\mathbf{A}^{c}_{t}=1,\mathbf{A}^{r}_{t}=0$), ($\mathbf{A}^{c}_{t}=0,\mathbf{A}^{r}_{t}=1$), ($\mathbf{A}^{c}_{t}=1,\mathbf{A}^{r}_{t}=1$)}. For $\tau$-step ahead prediction, we simulate synthetic tumor growth trajectories under single-sliding treatment (i.e., shift the treatment over a window) [bica2020estimating, melnychuk2022causal] and random trajectories (i.e., randomly assign treatments) settings [melnychuk2022causal]. Importantly, the ground-truth counterfactual trajectories are known. We limit the length of trajectories to a maximum of 60 time steps. For each setting, we generate trajectories with different amounts of time-varying confounding $\gamma\in[0,1,2,3,4]$, each with 10,000 trajectories for training, 1,000 for validation, and 1,000 for testing.
We follow the data simulation process and experimental setup as described in Appendix J and GitHub repository of the original Causal Transformer publication [melnychuk2022causal].
C.4 Semi-synthetic patient trajectories
MIMIC-III-CF is a semi-synthetic dataset based on patient data from real-world intensive care units [wang2020mimic, johnson2016mimic]. The data are aggregated at hourly levels, with forward and backward filling for missing values and standard normalization of the continuous time-varying features [wang2020mimic, johnson2016mimic, bica2020estimating, melnychuk2022causal]. Patients have 25 different vital signs as time-varying covariates and 3 static covariates (gender, ethnicity, age). Untreated trajectories of outcomes are first simulated under endogenous and exogenous dependencies, and then treatments are sequentially applied [melnychuk2022causal]. There are 3 synthetic binary treatments and 2 synthetic outcomes [melnychuk2022causal]. Importantly, the ground-truth counterfactual trajectories are known. We limit the length of trajectories to a maximum of 60 time steps. We generate 1,000 patients into train, validation, and test subsets via a 60%, 20%, and 20% split. For $\tau$-step ahead prediction with $\tau_{\text{max}}=10$, we sample 10 random trajectories for each patient per time step.
We follow the data simulation process and experimental setup as described in Appendix K and GitHub repository of the original Causal Transformer publication [melnychuk2022causal].
Appendix D Implementation Details
@@
We provide code and instructions in our GitHub repository to implement Clef, baselines, and ablations: https://github.com/mims-harvard/CLEF. To implement baselines, we follow the authors’ recommendations on model design and hyperparameter selection from the original publications. We do not share data or model weights that may contain sensitive patient information.
D.1 Baselines for Conditional Sequence Generation
We benchmark Clef against the state-of-the-art conditional sequence generation setup with different sequential data encoders (Figure 2): Transformer [vaswani2017attention, narasimhan2024time, jing2024towards, zhang2023survey] and xLSTM [beck2024xlstm]. We further compare against a state-of-the-art time series foundation model, MOMENT [goswami2024moment], as the sequence encoder. Specifically, we finetune an adapter for the $1024$-dimensional embeddings from the frozen MOMENT-1-large embedding model.
D.2 Clef Extensions and Baselines for Counterfactual Outcomes Estimation
Due to its versatility, Clef can be leveraged by state-of-the-art machine learning models designed to estimate counterfactual outcomes [bica2020estimating, melnychuk2022causal]. Counterfactual Recurrent Network (CRN) [bica2020estimating] and Causal Transformer (CT) [melnychuk2022causal] demonstrate state-of-the-art performance in the established benchmarks [bica2020estimating, melnychuk2022causal]. To implement Clef-CRN and Clef-CT, the GELU activation layer from the concept encoder $E$ (Equation 2) and the concept decoder $G$ (Equation 3) of Clef are appended to outcome predictor network (denoted as $G_{Y}$ where $Y$ is the outcome of the given treatment in the original publications [bica2020estimating, melnychuk2022causal]) of CRN and CT. Following the original CRN and CT publications, we minimize the factual outcome loss (i.e., output of $G_{Y}$) via mean squared error (MSE) [bica2020estimating, melnychuk2022causal].
We evaluate Clef against their non-Clef counterparts with and without balancing loss functions (i.e., gradient reversal (GR) [bica2020estimating], counterfactual domain confusion (CDC) loss [melnychuk2022causal]). This results in 5 distinct state-of-the-art baselines: CRN with GR loss (i.e., original CRN implementation) [bica2020estimating]; CRN with CDC loss [melnychuk2022causal]; CRN without balancing loss [melnychuk2022causal]; CT with CDC loss (i.e., original CT implementation) [melnychuk2022causal]; and CT without balancing loss [melnychuk2022causal].
D.3 Model Training
Models are trained on a single NVIDIA A100 or H100 GPU. All models have comparable number of parameters as their Clef-based counterparts.
D.4 Hyperparameter Sweep
@@
The selection of hyperparameters for the (conditional sequence generation) models trained from scratch are: dropout rate $\in[0.3,0.4,0.5,0.6]$, learning rate $\in[0.001,0.0001,0.00001]$, and number of layers (or blocks in xLSTM) $\in[4,8]$. Because the number of heads must be divisible by the number of features, the number of heads for eICU (18 lab tests) $\in[2,3,6,9]$ and for others $\in[4,8]$. For xLSTM, the additional hyperparameters are: 1D-convolution kernel size $\in[4,5,6]$ and QVK projection layer block size $\in[4,8]$.
D.5 Best Hyperparameters
@@
MIMIC-IV dataset. The best hyperparameters for the (conditional sequence generation) models trained on the MIMIC-IV dataset are: dropout rate $=0.6$, learning rate $=0.0001$, number of layers (blocks in xLSTM) $=8$, and number of heads $=4$. For xLSTM models, 1D-convolution kernel size $=4$ and QVK projection layer block size $=4$. For Clef models, the number of FNN in the concept encoder $=1$ (Appendix Figures 8-9).
@@
eICU dataset. The best hyperparameters for the (conditional sequence generation) models trained on the eICU dataset are: dropout rate $=0.6$, learning rate $=0.0001$, number of layers (blocks in xLSTM) $=8$, and number of heads $=6$. For xLSTM models, the number of heads $=2$, 1D-convolution kernel size $=4$ and QVK projection layer block size $=4$. For Clef models, the number of FNN in the concept encoder $=1$ (Appendix Figures 8-9).
@@
Waddington-OT (WOT) dataset. The best hyperparameters for the (conditional sequence generation) models trained on the WOT dataset are: dropout rate $=0.6$, learning rate $=0.00001$, number of layers (or blocks in xLSTM) $=4$, number of heads $=8$. For xLSTM models, 1D-convolution kernel size $=4$ and QVK projection layer block size $=8$. For Clef models, the number of FNN in the concept encoder $=0$ (Appendix Figures 8-9).
Synthetic tumor growth and semi-synthetic patient trajectories datasets. For the counterfactual prediction models (including Clef and non-Clef models), we follow the best hyperparameters as reported in the original publications of CT [melnychuk2022causal] and CRN [bica2020estimating].
Appendix E Additional Figures and Tables
@@
Figure 8: Benchmarking the performance of Clef, baselines, and ablation models on (a) immediate and (b) delayed sequence editing (Appendix C.1-C.2). Performance is measured by MAE (lower is better). Models are trained on 3 seeds using a standard cell- or patient-centric random split; error bars show 95% CI. Not shown for visualization purposes are the performances of VAR models on eICU and MIMIC-IV datasets: on immediate sequence editing, MAE for eICU and MIMIC-IV are $55982.74$ and $886.05$, respectively; on delayed sequence editing, MAE for eICU and MIMIC-IV are $3.02\times 10^{39}$ and $8.62\times 10^{23}$, respectively.
@@
Figure 9: Benchmarking the performance of Clef, baselines, and ablation models on (a) immediate and (b) delayed sequence editing (Appendix C.1-C.2). Performance is measured by RMSE (lower is better). Models are trained on 3 seeds using a standard cell- or patient-centric random split; error bars show 95% CI. Not shown for visualization purposes are the performances of VAR models on eICU and MIMIC-IV datasets: on immediate sequence editing, MAE for eICU and MIMIC-IV are $135003.67$ and $1793.23$, respectively; on delayed sequence editing, MAE for eICU and MIMIC-IV are $5.84\times 10^{39}$ and $1.59\times 10^{24}$, respectively.
@@
Figure 10: Cross-split overlap (CSO) as a function of SPECTRA parameter (SP) for eICU and MIMIC-IV datasets (Appendix C.2). CSO is defined as the number of samples in the test set that are similar to at least one sample in the train set. SP is an internal parameter used by SPECTRA to control the CSO of generated data splits. CSO decreases as SP increases. These data splits are used to evaluate conditional sequence generation models’ generalizability to unseen patient trajectories.
@@
Table 2: Generalizability of Clef, baselines, and ablations on eICU and MIMIC-IV datasets (Appendix C.2) in immediate and delayed sequencing. Performance is measured by the area under the spectral performance curve (AUSPC) for MAE (Appendix Figure 11) or RMSE (Appendix Figure 12). Smaller AUSPC values indicate better performance. Models are trained on 3 seeds; standard deviation is reported.
@@
+ Clef $15.16\pm 1.09$ $32.95\pm 2.47$ $14.36\pm 1.07$ $34.27\pm 2.12$ $32.79\pm 1.41$ $57.76\pm 3.39$ $35.65\pm 1.73$ $65.10\pm 4.43$
+ Clef + FFN $10.99\pm 0.31$ $27.57\pm 0.27$ $9.25\pm 0.60$ $27.69\pm 0.22$ $21.35\pm 3.16$ $36.92\pm 5.46$ $23.83\pm 3.26$ $44.11\pm 5.83$
@@
+ Clef $16.73\pm 2.16$ $35.43\pm 6.01$ $15.32\pm 2.10$ $34.68\pm 7.09$ $32.06\pm 1.13$ $53.42\pm 2.18$ $33.88\pm 1.98$ $57.73\pm 3.63$
+ Clef + FFN $11.35\pm 0.11$ $28.09\pm 0.08$ $9.04\pm 0.18$ $26.21\pm 0.48$ $21.04\pm 2.32$ $37.50\pm 4.60$ $22.63\pm 2.61$ $42.12\pm 5.03$
@@
+ Clef $47.69\pm 0.33$ $82.18\pm 0.34$ $40.10\pm 0.44$ $72.70\pm 0.46$ $44.01\pm 0.35$ $73.83\pm 0.63$ $46.88\pm 0.38$ $81.20\pm 1.27$
+ Clef + FFN $47.56\pm 1.60$ $82.81\pm 2.88$ $39.91\pm 1.65$ $72.54\pm 3.20$ $42.92\pm 0.52$ $70.72\pm 1.96$ $45.75\pm 0.65$ $77.35\pm 2.77$
@@
Figure 11: Generalizability of Clef, baselines, and ablation models on (a) eICU and (b) MIMIC-IV patient datasets (Appendix C.2) in immediate and delayed sequence editing. Performance is measured by MAE (lower is better). Models are trained on 3 seeds; error bars show 95% CI. As the SPECTRA parameter increases, the train/test split similarity decreases (Appendix Figure 10). The area under the spectral performance curve (AUSPC) evaluation is in Appendix Table 2.
@@
Figure 12: Generalizability of Clef, baselines, and ablation models on (a) eICU and (b) MIMIC-IV patient datasets (Appendix C.2) in immediate and delayed sequence editing. Performance is measured by RMSE (lower is better). Models are trained on 3 seeds; error bars show 95% CI. As the SPECTRA parameter increases, the train/test split similarity decreases (Appendix Figure 10). The area under the spectral performance curve (AUSPC) evaluation is in Appendix Table 2.
Figure 13: Counterfactual $\tau$-step ahead prediction on tumor growth (random trajectories setting) with different amounts of time-varying confounding $\gamma$ (Appendix C.3). GR refers to Gradient Reversal loss [bica2020estimating]; CDC refers to Counterfactual Domain Confusion loss [melnychuk2022causal]; BL refers to Balancing Loss (i.e., GR or CDC). Models are trained on 5 seeds; error bars show 95% CI.
Figure 14: Counterfactual $\tau$-step ahead prediction on tumor growth (single-sliding treatment) with different amounts of time-varying confounding $\gamma$ (Appendix C.3). GR refers to Gradient Reversal loss [bica2020estimating]; CDC refers to Counterfactual Domain Confusion loss [melnychuk2022causal]; BL refers to Balancing Loss (i.e., GR or CDC). Models are trained on 5 seeds; error bars show 95% CI.
Figure 15: Counterfactual $\tau$-step ahead prediction on tumor growth (random trajectories setting) with different amounts of time-varying confounding $\gamma$ (Appendix C.3). GR refers to Gradient Reversal loss [bica2020estimating]; CDC refers to Counterfactual Domain Confusion loss [melnychuk2022causal]; BL refers to Balancing Loss (i.e., GR or CDC). Models are trained on 5 seeds; error bars show 95% CI.
Figure 16: Counterfactual $\tau$-step ahead prediction on semi-synthetic patient trajectories (Appendix C.4). GR refers to Gradient Reversal loss [bica2020estimating]; CDC refers to Counterfactual Domain Confusion loss [melnychuk2022causal]; BL refers to Balancing Loss (i.e., GR or CDC). Models are trained on 5 seeds; error bars show 95% CI.
@@
Figure 17: Benchmarking the performance of Clef, baselines, and ablation models on zero-shot (a) immediate and (b) delayed counterfactual generation of cellular developmental trajectories (Appendix C.1). Performance is measured by MAE (top row) and RMSE (bottom row). Models are trained on 3 seeds; error bars show 95% CI.